OpenStudy (anonymous):

how do i estimate a 51%of 62 and write it as a fraction

OpenStudy (anonymous):

( 51 / 100 ) * 62 = 1581/50

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

can u help me on another

OpenStudy (anonymous):

Or if you don't like fractions- (.51)*(62)=31.62

OpenStudy (anonymous):

no we have to do fractions

OpenStudy (anonymous):

what is 39%of 42 as a fraction

OpenStudy (anonymous):

estimated to

OpenStudy (anonymous):

Same way. (39/100)*42=16.38

OpenStudy (anonymous):

Same way. (39/100) * 42 = 819/50

OpenStudy (anonymous):

what is 78%of148 estimated and wrote as a fraction

OpenStudy (anonymous):

when you say (numer)%<--- it means it's (number)/(100) and when you say (of)it means mulitplication :)

OpenStudy (anonymous):

(78/100) * 148 =2886/25

OpenStudy (anonymous):

number* do you understand what's going on now ^_^?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol, alright

OpenStudy (anonymous):

\[(number)% \rightarrow \frac{number}{100}\]

OpenStudy (anonymous):

what is 34% of 99 estimated and wrote as a fraction

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

dumb-dude--- rewrite it the same way we've explained the last problems.

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

dum , 34% is a number, and it's 34/100 in fraction

OpenStudy (anonymous):

It's (the percent) times the number.

OpenStudy (anonymous):

like I've said before: (number)% = (number)/100 anything that has the %<- sign next to it is always over 100 ^_^

OpenStudy (anonymous):

exactly like what Murph said ^_^

OpenStudy (anonymous):

oh ok i get it so 34/100 and then times those numbers right

OpenStudy (anonymous):

yep ^_^

OpenStudy (anonymous):

whenever you have "of" it means multiply :)

OpenStudy (anonymous):

ok guys thanks for ur help i love this website

OpenStudy (anonymous):

np ^_^ and good luck :)

OpenStudy (anonymous):

34/100 * 99/1 and just multiply straight across. Lastly, simplify the fraction you end up with.

OpenStudy (anonymous):

what does the star and 99 mean

OpenStudy (anonymous):

* means multiplty

OpenStudy (anonymous):

star = x( multiplication)

OpenStudy (anonymous):

99 was the number you gave us

OpenStudy (anonymous):

yes so u take 34 divided by 1oo and then u multiply 99

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

well done ~ :)

OpenStudy (anonymous):

Keep doing it over and over. Repitition is the best way to learn. Good luck!